For this analysis, you will want to refer to the section in chapter 7 of the Baye/Prince text on the computation of the Herfindahl-Hirschman Index and the section on firm “conduct” dealing with “integration and merger activity,” particularly “horizontal integration.” You’ll also find a bit more commentary on mergers and antitrust legislation near the beginning of chapter 14, “A Manager’s Guide to Government in the Marketplace.”

Industry data is from the IBISWorld data base, accessible to JBU students through the JBU Library’s online resources. According to IBISWorld: “Pharmaceutical manufacturers develop prescription and over-the-counter products that are used to prevent or treat illnesses in humans or animals. Brand name drugs and medication have patent protection. This industry does not include nutritional supplement or cosmetic product manufacturers.”

Let’s do a little industry analysis; the numbers here a generally reflective of market conditions in recent years, with minor adjustments, but this exercise should still be considered “hypothetical.” Suppose annual industry sales are $188.4 billion. In this industry there are six leading companies, several smaller significant competitors, and a number of even smaller niche firms. The top companies and their revenue data are:

Company

Revenue

Pfizer Inc.

$ 22.2 billion

AbbVie Inc.

$ 21.5 billion

Johnson & Johnson

$ 21.5 billion

Merck & Co. Inc.

$ 19.4 billion

Amgen Inc.

$ 18.3 billion

Bristol-Myers Squibb

$ 13.0 billion

Eli Lilly & Co.

$ 9.2 billion

Sanofi SA

$ 9.2 billion

GlaxoSmithKline PLC

$ 6.4 billion

Other smaller firms combined

$ 47.7 billion

Industry Total

$188.4 billion

Suppose that Amgen announced it intended to merge with Eli Lilly. What would be the new industry HHI? Given the policies stated in the “Horizontal Merger Guidelines” as discussed in the text, would the merger most likely be challenged by the DOJ/FTC as anticompetitive?
Answer #1

Solution

Company

Revenue

Pfizer Inc.

$ 22.2 billion

AbbVie Inc.

$ 21.5 billion

Johnson & Johnson

$ 21.5 billion

Merck & Co. Inc.

$ 19.4 billion

Amgen Inc.

$ 18.3 billion

Bristol-Myers Squibb

$ 13.0 billion

Eli Lilly & Co.

$ 9.2 billion

Sanofi SA

$ 9.2 billion

GlaxoSmithKline PLC

$ 6.4 billion

Other smaller firms combined

$ 47.7 billion

Industry Total

$188.4 billion

The market shares of all the above players are:

Pfizer =  22.2 / 188.40 i.e.,11.78 %

Abbvive= 21.50 / 188.40 i.e.,11.41%

Johnson & Johnson = 21.50 / 188.40 i.e.,11.41%

Merck & Co. Inc. = 19.40 / 188.40 i.e., 10.29%

Amgen = 18.30 / 188.40 i.e.,9.71%

Bristol-Myers Squibb = 13.0 / 188.40 i.e., 6.90 %

Eli Lilly & Co. = 9.20 / 188.40 i.e., 4.88 %

Sanofi SA = 9.20 / 188.40 i.e., 4.88 %

GlaxoSmithKline PLC = 6.40 / 188.40 i.e., 3.39 %

Other smaller firms combined = 47.70 / 188.40 i.e., 25.31 %

HHI of the industry is : (11.78)2 + (11.41)2 + (11.41)2 +(10.29)2​​​​​​​ +(9.71)2​​​​​​​ +(6.90)2​​​​​​​ + (4.88)2​​​​​​​ + (4.88)2​​​​​​​ + (3.39)2​​​​​​​ +(25.31)2​​​​​​​

= 138.76 + 130.18 + 130.18 +105.88 + 94.28 + 47.61 + 23.81 + 23.81 + 11.49 + 645.66

= 1351.66

Since a  HHI index of less than 1500 indicates that the market is competitive,the above market is competitive i.e., the market place is not at all concentrated

Suppose Amgen announced a merger with Eli Lilly then the new HHI would be :

= (11.78)2 + (11.41)2 + (11.41)2​​​​​​​ +(10.29)2​​​​​​​ +(9.71 +4.88)2​​​​​​​ +(6.90)2​​​​​​​ + (4.88)2​​​​​​​ + (3.39)2​​​​​​​ +(25.31)2​​​​​​​

= 138.76 + 130.18 + 130.18 +105.88 + 212.86 + 47.61 + 23.81 + 11.49 + 645.66

= 1446.43

So,still the market remains to be competitive .

Here the change in the HHI after the merger (HHI Post)is : 1446.43 - 1351.66 i.e., 94.77

According to the DOJ / FTC Guidelines,since the HHI post the merger is in between 1000 & 1800 ; and that the change in the HHI after the merger is less than 100,this merger is unlikely to be challenged as anti-competitive.

Hope this solution helps !! Please give a " Thumbs Up" rating for this solution !!

​​​​​​​